OpenStudy (anonymous):

| 3cis(3pi/4) | preform the following operation and state the answer in trig form

OpenStudy (anonymous):

poop

OpenStudy (sirm3d):

\[3cis(3\pi/4)\] is a shorthand for \[3\cos(3\pi/4)+3i\sin(3\pi/4)\]

OpenStudy (anonymous):

yes i know